WARNING: This web is not used anymore. Please use PDBService.MinuteS18January2007 instead!
 

Minutes 18 January 07

  • Phone:
    • TRIUMF: Denice
    • GridKA: Andreas
    • FNAL: Eric
    • ATLAS: Shasha
    • SARA: Alexander
    • BNL: Iris
    • NDGF: Olli
    • RAL: Gordon
    • LHCB: Marco

  • CERN:
    • 3D: Dirk, Zbigniew, Marcin, Luis, Gianni, Eva
    • ATLAS: Gancho, Florbela
    • COOL: Andrea

  • General announcements:
    • workshop next week 22 - 26 January: phone conference will be available
    • next dba meeting - 20, 21 March - SARA

  • Sites update:
    • RAL (Gordon): email list for the streams monitoring tool provided
    • NDGF (Olli):
      • machine is up and running
      • Oracle ASM installation ongoing - database installation to be done after the workshop
      • looking at how to adquire the required hw (3 nodes)
    • BNL (Iris):
      • memory has been increased to 2GB
      • network problems being investigated
    • Triumf (Denice):
      • working on redundancy for fiber chanel
      • oem agents being installed
    • SARA (Alexander):
      • working with Florbela and Gancho configuring the streams tests
      • problems with the configuration of the listener

  • ATLAS Streams tests:
    • Streams setup to RAL for Richard's tests
    • Tier 1 sites: is it possible to upgrade to 10.2.0.3 early february?
    • Quarterly security patch: the impact on 10.2.0.2 is minimal but the sites should review the documentation to determine which are their vulnerabilities.
    • The condition challenge will be end february

  • LHCB replication status:
    • Marco asks for an oem account - should contact Dirk Geppert
    • CERN to GridKA, RAL and IN2P3 being tested without problems so far
    • CNAF to be added soon - already contacted
    • Jobs will be submitted through the grid to all the destination sites
    • Performance and functionality tests

  • CMS replication
    • redo log generation on the source database is lower than on the destination database
    • factor 10 observed
    • additional redo seems to be related to the spill over to disk
    • Oracle is involved in order to fix this
    • redo monitoring has being added to the Streams monitoring tool (Zbigniew)
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-02-01 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback